3. 168 points E2-16 Calculating Actual and Applied Manufacturing Overhead Costs and Over- or Underapplied Overhead Costs [LO 2-3, 2-4, 2-5. 2-6) Verizox Company uses a job order cost system with menufacturing overhead appied to products based on direct labor hours At the beginning of the most recent year, the company estimated its manufacturing overhead cost at $183,080. Estimated diract labor cost was $471630 for 19,900 hours Actual costs for the most recent month are summarized hare Hem Descrietion Direct iabor i1.900 hours) indroct cosfs Total Cost 47,164 2400 3.380 Indirect materials Fadory rent Facdory supenvision Factory depreciation Factiory jaritorial wok Facdory insurance Gerrd and administradive saaries Beling sspenses 4,710 5,600 1,180 4,140 5.300 Required 1. Calcuale the predstermined owerhead rate (Roand your ans war to 2 decimal places.) Per DL Hour 2. Caioulate the amount af applied manufacturing averhead 3. Cakcuate achual manutacturing overhead costs 4. Compute over-or underappiod o.erhead, tinput the amount as poitive value ,
Answer #1
1. Predetermined overhead rate = $183080 / 19900 = $9.20 per DL hour
2. Applied manufacturing overhead = Predetermined overhead rate × Actual value of allocation base
Applied manufacturing overhead = $9.20 × 1,800 actual direct labor hours = $16560
3. Actual Manufacturing Overhead Costs = Indirect Labor + Indirect Material + Factory Rent + Factory Supervision + Factory Depreciation + Factory Janitorial Work + Factory Insurance = 2570+3310+3250+4840+5690+1140+1780
$ 22,580.00
4. Overhead = Actual Manufacturing Overhead - Applied manufacturing overhead = $22580-16560 = $6020(Overapplied)